# Real-world models for multiple term structures: a unifying HJM semimartingale framework

### v2

@@ 1. Introduction
    We propose an abstract parameterization of a market with multiple term structures and study it from the viewpoint of large financial markets consisting of uncountably many assets as in [18]. We work under the real-world probability and derive a version of the fundamental theorem of asset pricing which characterizes market viability (as introduced in [38] for finite-dimensional semimartingale models) for large financial markets under an infinite time horizon.
@@
    We develop a semimartingale HJM framework for multiple term structures under the real-world probability imposing only mild regularity conditions. We provide a complete characterization of the set of local martingale deflators, whose existence ensures market viability, and show that it can be explicitly characterized in terms of drift restrictions that generalize the HJM condition of [33]. Moreover, we derive a simple condition which ensures ordered term structures.
@@
    We prove a new existence and uniqueness theorem for a class of semilinear stochastic partial differential equations (SPDEs) with random locally Lipschitz coefficients, driven by a Brownian motion and a Poisson random measure. By relying on this general result, we establish the well-posedness of the HJM semimartingale model by proving existence and uniqueness of solutions to the associated SPDE. Besides extending the results of [25] to a multi-dimensional setting, we substantially weaken the technical requirements, thereby covering new classes of models that cannot be treated by the existing theory. Moreover, we analyze the invariance properties of the SPDE and provide conditions for the existence of finite-dimensional (affine) realizations.
@@ 1.1.1. Foreign exchange markets
A first example of multiple term structures arises from yield curves in international financial markets. Similarly to [37], this example will inspire our general parameterization of multiple term structures. We consider a domestic economy associated to a reference currency and a family $I=\{1,\ldots,m\}$ of foreign economies, each of them associated to a distinct foreign currency. For each $i\in I$, the value at time $t$ of one unit of the $i$-th foreign currency in units of the domestic currency is given by the spot exchange rate $S^{i}_{t}$, while the $i$-th yield curve is specified by foreign ZCB prices $B^{i}(t,T)$, representing the value at time $t$ in units of the $i$-th foreign currency of one unit of the same currency delivered at time $T\geq t$. From the perspective of a domestic investor, the value of the $i$-th foreign ZCB at time $t$ is $S^{i}_{t}B^{i}(t,T)$, for each $i\in I$. Assuming that international trading is allowed, foreign ZCBs constitute risky assets for domestic investors due to currency risk, unlike domestic ZCBs. This setup naturally gives rise to $m+1$ term structures, corresponding to the domestic yield curve and the $m$ foreign yield curves, which must coexist in an arbitrage-free way. HJM models for FX markets have been first introduced in [1] and later analyzed in a semimartingale setup in [42].
@@ 1.1.2. Interbank interest rates
Multiple term structures arise in interest rate markets when considering benchmark rates indexed by a family $I=\{\delta_{1},\ldots,\delta_{m}\}$ of ordered tenors (see [32] for an overview on multi-curve interest rate models). The most basic contract referencing a benchmark rate is a single-period swap (forward rate agreement), in which at maturity $T+\delta_{i}$ the benchmark rate $L_{T}(\delta_{i})$ fixed at time $T$ for tenor $\delta_{i}$ is exchanged against a fixed rate $K$. It has been shown in [29] that the value of this contract at time $t\leq T$ is given by $S^{i}_{t}B^{i}(t,T)-(1+\delta_{i}K)B^{0}(t,T+\delta_{i})$, where $S^{i}_{t}:=(1+\delta L_{t}(\delta_{i}))B^{0}_{t}(t+\delta_{i})$ is the multiplicative spread at time $t$ between the $\delta_{i}$-tenor rate and the simply compounded risk-free rate over the interval $[t,t+\delta_{i}]$ and $B^{i}(t,T)$ is a term structure factor satisfying $B^{i}(T,T)=1$, for all $i=1,\ldots,m$ and $T\geq 0$. The tenor dependence arises due to the distinct impact of credit, funding, and liquidity risks over different tenors in the interbank market. Since longer tenors are associated to greater risks, the quantity $S^{i}_{t}B^{i}(t,T)$ is generally increasing with respect to $i$. The existence of swaps of different maturities referencing benchmark rates of different tenors, together with the risk-free term structure, results in a market setup with $m+1$ term structures, which must coexist in an arbitrage-free way. Multi-curve interest rate models have been studied in a semimartingale setup in [16, 17, 29].
@@ 1.1.4. Longevity bonds
Multiple term structures also arise in insurance markets where longevity bonds are traded, i.e., zero-coupon bonds whose payoff depends on the value of a specified survivor index. We consider a collection of $m$ survivor indices, each corresponding to a different age group. For each $i=1,\ldots,m$, let $S^{i}_{t}$ denote the $i$-th survivor index at time $t$, defined as the proportion of individuals alive at time $t$ from the initial cohort belonging to the $i$-th age group (survival ratio). A longevity bond maturing at $T$ linked to the $i$-th survivor index has payoff $S^{i}_{T}$ at time $T$. The market value of this bond at time $t\leq T$ can be expressed as $S^{i}_{t}B^{i}(t,T)$, with $B^{i}(t,T)$ satisfying the terminal condition $B^{i}(T,T)=1$, for all $T\geq 0$. Together with the risk-free term structure, this setup gives rise to $m+1$ term structures coexisting in the same market. Moreover, if the age groups are ordered increasingly, then $S^{i}_{t}B^{i}(t,T)$ should be decreasing in $i$, as longevity bonds linked to older cohorts should trade at lower prices. Adopting this parameterization of longevity bonds, a HJM model for the term structure of longevity bonds has been first developed in [3] (see also [53] for a more general HJM formulation).
@@ 1.1.5. Energy forward contracts
In energy markets, swap contracts for the delivery of energy over a specified future time interval $[T,T+\delta]$ are widely traded (this is for instance the case of flow forward contracts on electricity, gas and temperature), where $\delta$ is the delivery length. Let $I$ be an index set corresponding to all possible delivery lengths. For each $i\in I$, let $S^{i}_{t}$ denote the spot price of a swap contract with immediate delivery over the interval $[t,t+\delta_{i}]$, for all $t\geq 0$. The market value at time $t$ of a swap contract delivering energy over $[T,T+\delta_{i}]$, with $T\geq t$, can then be expressed as $S^{i}_{t}B^{i}(t,T)$, with $B^{i}(t,T)$ representing a forward adjustment factor for maturity $T$ associated with delivery length $\delta_{i}$, satisfying $B^{i}(T,T)=1$. This setup generates multiple term structures which must coexist in an arbitrage-free way, since swap contracts for different delivery lengths are traded in the same market. If the delivery lengths are increasingly ordered, then also $S^{i}_{t}B^{i}(t,T)$ should be increasing in $i$, thereby implying a monotonicity property among the term structures. A HJM model for swap contracts with different delivery lengths has been proposed in [4] and then generalized in [5], [6, Chapter 6] and [7].
@@ 2.2. Market viability
In this section, we derive a version of the fundamental theorem of asset pricing that is applicable to an abstract financial market with multiple term structures as introduced above. We adopt a mild notion of market viability (Definition 2.2) corresponding to the no unbounded profit with bounded risk (NUPBR) condition of [38]. [^1] We refer to [29] for a version of the fundamental theorem of asset pricing for multi-curve interest rate models based on the stronger notion of no asymptotic free lunch with vanishing risk (see [18]). The abstract financial market, as introduced in Definition 2.1, is a large financial market with uncountably many assets. In order to define and characterize market viability, we adopt the approach of [18] (see in particular Example 2.2 therein [^2] We point out that the index set $I_{0}\times\mathbb{R}_{+}$ identifying the assets of the financial market of Definition 2.1 can be mapped in a bijective way onto a subset of $\mathbb{R}_{+}$, therefore satisfying the requirements of a parameter space as considered in [18].), considering wealth processes that are limits in the semimartingale topology of admissible portfolios involving a finite number of arbitrarily chosen assets. More precisely, for each $n\in\mathbb{N}$, we define
@@
$\mathcal{X}^{A}_{1}:=\bigl{\{}H\cdot X^{A}:H\in L(X^{A})\text{ and }H\cdot X^{A}\geq-1\bigr{\}},$
@@
Definition 2.2 extends [18, Definition 4.1] to an infinite time horizon by requiring NUPBR to hold over every arbitrary finite time horizon. This extension is analogous to the one first adopted in [39] in the context of a finite-dimensional market. We remark that requiring NUPBR on $[0,T]$, for all $T>0$, is weaker than having NUPBR on $[0,+\infty)$ as first considered in [38] (see [2, Remark 5.3]).
@@
A strictly positive càdlàg process $Z$ with $Z_{0}\leq 1$ is a supermartingale deflator if $Z(1+X)$ is a supermartingale, for every $X\in\mathcal{X}_{1}$. A strictly positive local martingale $Z$ with $Z_{0}=1$ is a local martingale deflator (LMD) if $ZX^{(i,T)}$ is a local martingale, for every $(i,T)\in I_{0}\times\mathbb{R}_{+}$.
The next theorem is a version of the fundamental theorem of asset pricing based on the above notion of NUPBR. This result extends [18, Theorem A.1] in two directions: first, by considering an infinite time horizon; second, by proving the existence of a supermartingale deflator which is the reciprocal of a wealth process (numéraire portfolio, see [38]). While the extension is relatively straightforward, relying on existing results such as [40, Theorem 1.7], we give a detailed proof in Appendix A, as we could not find in the literature a formulation that is directly applicable to our setting. We note that, in the present setup, the process $1+\widehat{X}$ appearing in Theorem 2.5 corresponds to the wealth process of the growth-optimal portfolio, a central object in the benchmark approach to finance (see, e.g., [46]).
@@
For the financial market of Definition 2.1, NUPBR holds if and only if there exists an element $\widehat{X}\in\mathcal{X}_{1}$ with $\widehat{X}>-1$ such that $1/(1+\widehat{X})$ is a supermartingale deflator. Moreover, every local martingale deflator is a supermartingale deflator.
We point out that, at the present level of generality, NUPBR does not imply the existence of LMDs. Indeed, an explicit example in [18] shows that in large financial markets market viability (in the sense of NUPBR, but also in the stronger sense of no asymptotic free lunch with vanishing risk) does only ensure the existence of supermartingale deflators, and not necessarily of LMDs. [^3] In the specific case of continuous-path processes NUPBR always implies the existence of an LMD, as shown in [41]. However, given a concrete model, it is usually difficult to characterize supermartingale deflators, since their definition is based on the whole set $\mathcal{X}_{1}$ and does not give information on the basic assets included in Definition 2.1. On the contrary, LMDs can be directly described in terms of the characteristics of the basic assets, as we are going to show in Section 3. Whenever it is nonempty, we denote by $\mathcal{D}$ the set of all LMDs.
@@ 3.1. Probabilistic setup
Let the filtered probability space $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ support a $d$-dimensional Brownian motion $W$ and an integer-valued random measure $\mu({\rm d}t,{\rm d}x)$ on +×$\mathbb{R}_{+}\times E$ with compensator $\nu({\rm d}t,{\rm d}x)=F_{t}({\rm d}x){\rm d}t$, where $E$ is a Polish space with its Borel sigma-field $\mathcal{B}(E)$ and $F_{t}({\rm d}x)$ is a kernel from $(\Omega\times\mathbb{R}_{+},\mathcal{P})$ into $(E,\mathcal{B}(E))$, with $\mathcal{P}$ denoting the predictable sigma-field on $\Omega\times\mathbb{R}_{+}$. The compensated random measure is denoted by $\tilde{\mu}({\rm d}t,{\rm d}x):=\mu({\rm d}t,{\rm d}x)-\nu({\rm d}t,{\rm d}x)$. We denote by $L^{2}_{\rm loc}(W)$ the set of all progressively measurable $\mathbb{R}^{d}$-valued processes $\theta=(\theta_{t})_{t\geq 0}$ such that $\int_{0}^{T}\|\theta_{t}\|^{2}{\rm d}t<+\infty$ a.s., for all $T>0$, and by $G_{\rm loc}(\mu)$ the set of all $\mathcal{P}\otimes\mathcal{B}(E)$-measurable functions $\varphi:\Omega\times\mathbb{R}_{+}\times E\to\mathbb{R}$ such that $\int_{0}^{T}\int_{E}((\varphi^{i}_{t}(x))^{2}\wedge|\varphi^{i}_{t}(x)|)F_{t}({\rm d}x){\rm d}t<+\infty$ a.s., for all $T>0$ (compare with [35, Theorem II.1.33-c]). We refer the reader to [35] for all unexplained notions of stochastic calculus.
@@
    $\int_{0}^{T}\int_{E}\int_{0}^{T}(\gamma^{i}(s,u,x))^{2}{\rm d}u\,F_{s}({\rm d}x){\rm d}s<+\infty,\qquad\text{ for all }T>0.$
Assumption 3.1 implies that the integrals appearing in the forward rate equation (3.2) are well-defined for a.e. $T$. In addition, the integrability requirements appearing in parts (ii)-(iv) of Assumption 3.1 ensure the applicability of ordinary and stochastic Fubini theorems, in the versions of [54, Theorem 2.2] for the Brownian motion $W$ and [10, Proposition A.2] for the compensated random measure $\tilde{\mu}$. By [54, Remark 2.1], the mild measurability requirement in parts (ii)-(iii) holds if the processes $\alpha^{i}$ and $\beta^{i}$ are $\mathrm{Prog}\otimes\mathcal{B}(\mathbb{R}_{+})$-measurable, with $\mathrm{Prog}$ denoting the progressive sigma-field on $\Omega\times\mathbb{R}_{+}$.
@@
(2) Some recent works (see for instance [29, 30]) have considered HJM models that do not satisfy quasi-left-continuity (i.e., the set $\{(\omega,t)\in\Omega\times\mathbb{R}_{+}:\nu(\omega;\{t\}\times E)>0\}$ is not evanescent). Theorem 3.4 below can be generalized to this situation with an analogous proof. However, we do not purse this generalization here since the SPDE analysis of Section 4 will necessitate quasi-left-continuity.
@@
[…]
@@
[…]
thus ensuring that the stochastic integral $\int_{0}^{\cdot}\int_{E}\bar{\gamma}^{i}(s,x,T)\tilde{\mu}({\rm d}s,{\rm d}x)$ is well-defined as a local martingale, see [35, Theorem II.1.33]. The finiteness of the integral $\int_{0}^{\cdot}f^{i}(s,s){\rm d}s$ follows similarly as above under the validity of Assumption 3.1. An application of [35, Theorem II.8.10] to (3.3) gives the representation
@@
×$\times\exp\left(\int_{0}^{t}f^{i}(s,s){\rm d}s-\int_{0}^{t}\bar{\alpha}^{i}(s,T){\rm d}s+\frac{1}{2}\int_{0}^{t}\|\bar{\beta}^{i}(s,T)\|^{2}{\rm d}s\right).$
The result of the lemma then follows by an application of Yor’s formula (see, e.g., [35, formula II.8.19]), making use of (3.4) and recalling that $S^{i}=S^{i}_{0}\mathcal{E}(Z^{i})$, where $Z^{i}$ is given by (3.1). ∎

### v3

@@ 1. Introduction
    We introduce an abstract semimartingale parameterization for markets with multiple term structures and analyze it from the viewpoint of large financial markets with uncountably many assets, in the spirit of [18]. We work under the real-world probability and derive a version of the fundamental theorem of asset pricing which characterizes market viability (as introduced in [41] for finite-dimensional semimartingale models) for large financial markets under an infinite time horizon. Working under the real-world probability is practically relevant because viable market models may not admit a risk-neutral probability, while valuation and hedging problems can still be meaningfully posed and consistently solved, as shown for instance in [21, 32, 42, 51].
@@
    We develop a semimartingale HJM framework for multiple term structures under the real-world probability imposing only mild regularity conditions. We provide a complete characterization of the set of local martingale deflators, whose existence ensures market viability, and show that it can be explicitly characterized in terms of drift restrictions that generalize the classical HJM condition of [35]. We also derive a simple condition which ensures ordered term structures. As a special case of our general framework, we recover the risk-neutral setup and obtain risk-neutral HJM drift restrictions under conditions weaker than those found in the literature.
@@
    We prove a new existence and uniqueness theorem for a class of semilinear stochastic partial differential equations (SPDEs) with random locally Lipschitz coefficients, driven by a Brownian motion and a Poisson random measure. By relying on this general result, we establish the well-posedness of the HJM semimartingale dynamics by proving existence and uniqueness of solutions to the associated SPDE. Besides extending the results of [25] to a multi-dimensional setting, we substantially weaken the technical requirements, thereby covering new classes of models that cannot be treated by the existing theory, also in the risk-neutral setup, as shown by means of an explicit example. Finally, we analyze the invariance properties of the SPDE and provide conditions for the existence of finite-dimensional (affine) realizations.
@@ 1.1.1. Foreign exchange markets
A first example of multiple term structures arises from yield curves in international financial markets. Similarly to [39], this example will inspire our general parameterization of multiple term structures. We consider a domestic economy associated to a reference currency and a family $I=\{1,\ldots,m\}$ of foreign economies, each of them associated to a distinct foreign currency. For each $i\in I$, the value at time $t$ of one unit of the $i$-th foreign currency in units of the domestic currency is given by the spot exchange rate $S^{i}_{t}$, while the $i$-th yield curve is specified by foreign ZCB prices $B^{i}(t,T)$, representing the value at time $t$ in units of the $i$-th foreign currency of one unit of the same currency delivered at time $T\geq t$. From the perspective of a domestic investor, the value of the $i$-th foreign ZCB at time $t$ is $S^{i}_{t}B^{i}(t,T)$, for each $i\in I$. If international trading is allowed, foreign ZCBs constitute risky assets for domestic investors due to currency risk, unlike domestic ZCBs. This framework yields $m+1$ term structures, corresponding to the domestic yield curve and the $m$ foreign yield curves, which must coexist in an arbitrage-free way. HJM models for FX markets were first introduced in [1] and later analyzed in a semimartingale setup in [46]. Inflation-linked term structure models, as considered for instance in the seminal framework of [40], also naturally fit into this setting, since they involve the joint modeling of both nominal and real interest rate term structures.
@@ 1.1.2. Interbank interest rates
Multiple term structures arise in interest rate markets when considering benchmark rates indexed by a family $I=\{\delta_{1},\ldots,\delta_{m}\}$ of ordered tenors (see [34] for an overview of multi-curve interest rate models). The most basic contract referencing a benchmark rate is a single-period swap (forward rate agreement), in which at maturity $T+\delta_{i}$ the benchmark rate $L_{T}(\delta_{i})$ fixed at time $T$ for tenor $\delta_{i}$ is exchanged for a fixed rate $K$. It has been shown in [29] that the value of this contract at time $t\leq T$ equals $S^{i}_{t}B^{i}(t,T)-(1+\delta_{i}K)B^{0}(t,T+\delta_{i})$, where $S^{i}_{t}:=(1+\delta L_{t}(\delta_{i}))B^{0}_{t}(t+\delta_{i})$ is the multiplicative spread at time $t$ between the $\delta_{i}$-tenor rate and the simply compounded risk-free rate over $t,t+\delta_{i}]$ and $B^{i}(t,T)$ is a term structure factor with $B^{i}(T,T)=1$, for all $i=1,\ldots,m$ and $T\geq 0$. The tenor dependence reflects the differing impact of credit, funding, and liquidity risks across tenors in the interbank market. Since longer tenors involve greater risks, $S^{i}_{t}B^{i}(t,T)$ is generally increasing in $i$. The existence of swaps of different maturities referencing benchmark rates of different tenors, together with the risk-free term structure, leads to a market with $m+1$ term structures that must coexist in an arbitrage-free way. Multi-curve interest rate models have been studied in a semimartingale setup in [16, 17, 29]. Similarly, multiple term structures also arise in the cross-currency HJM framework of [33], where each term structure is indexed by both a base and a collateral currency.
@@ 1.1.4. Longevity bonds
Multiple term structures also arise in insurance markets where longevity bonds are traded, i.e., zero-coupon bonds whose payoff depends on the value of a specified survivor index. We consider a collection of $m$ survivor indices, each corresponding to a different age group. For each $i=1,\ldots,m$, let $S^{i}_{t}$ denote the $i$-th survivor index at time $t$, defined as the proportion of individuals alive at time $t$ from the initial cohort belonging to the $i$-th age group (survival ratio). A longevity bond maturing at $T$ linked to the $i$-th survivor index has payoff $S^{i}_{T}$ at time $T$. The market value of this bond at time $t\leq T$ can be expressed as $S^{i}_{t}B^{i}(t,T)$, with $B^{i}(t,T)$ satisfying the terminal condition $B^{i}(T,T)=1$, for all $T\geq 0$. Together with the risk-free term structure, this setup gives rise to $m+1$ term structures coexisting in the same market. Moreover, if the age groups are ordered increasingly, then $S^{i}_{t}B^{i}(t,T)$ should be decreasing in $i$, as longevity bonds linked to older cohorts should trade at lower prices. Adopting this parameterization of longevity bonds, a HJM model for the term structure of longevity bonds was first developed in [3] (see also [58] for a more general HJM formulation).
@@ 1.1.5. Energy forward contracts
In energy markets, swap contracts for the delivery of energy over a specified future time interval $T,T+\delta]$ are widely traded (this is for instance the case of flow forward contracts on electricity, gas and temperature), where $\delta$ is the delivery length. Let $I$ be an index set corresponding to all possible delivery lengths. For each $i\in I$, let $S^{i}_{t}$ denote the spot price of a swap contract with immediate delivery over the interval $t,t+\delta_{i}]$, for all $t\geq 0$. The market value at time $t$ of a swap contract delivering energy over $T,T+\delta_{i}]$, with $T\geq t$, can then be expressed as $S^{i}_{t}B^{i}(t,T)$, with $B^{i}(t,T)$ representing a forward adjustment factor for maturity $T$ associated with delivery length $\delta_{i}$, satisfying $B^{i}(T,T)=1$. This setup generates multiple term structures which must coexist in an arbitrage-free way, since swap contracts for different delivery lengths are traded in the same market. If the delivery lengths are increasingly ordered, then also $S^{i}_{t}B^{i}(t,T)$ should be increasing in $i$, thereby implying a monotonicity property among the term structures. A HJM model for swap contracts with different delivery lengths has been proposed in [4] and then generalized in [5], [6, Chapter 6] and [7].
@@ 2.2. Market viability
In this section, we derive a version of the fundamental theorem of asset pricing that is applicable to an abstract financial market with multiple term structures as introduced above. We adopt a mild notion of market viability (Definition 2.2) corresponding to the no unbounded profit with bounded risk (NUPBR) condition of [41]. [^1] We refer to [29] for a version of the fundamental theorem of asset pricing for multi-curve interest rate models based on the stronger notion of no asymptotic free lunch with vanishing risk introduced in [18]. The abstract financial market, as introduced in Definition 2.1, is a large financial market with uncountably many assets. In order to define and characterize market viability, we adopt the approach of [18] (see in particular Example 2.2 therein [^2] We point out that the index set $I_{0}\times\mathbb{R}_{+}$ identifying the assets of the financial market of Definition 2.1 can be mapped in a bijective way onto a subset of $\mathbb{R}_{+}$, therefore satisfying the requirements of a parameter space as considered in [18].), considering wealth processes that are limits in the semimartingale topology of admissible portfolios involving a finite number of arbitrarily chosen assets. More precisely, for each $n\in\mathbb{N}$, we define
@@
[…]
@@
Definition 2.2 extends [18, Definition 4.1] to an infinite time horizon by requiring NUPBR to hold over every arbitrary finite time horizon. This extension is analogous to the one first adopted in [43] in the context of a finite-dimensional market. We remark that requiring NUPBR on $[0,T]$, for all $T>0$, is weaker than having NUPBR on $0,+\infty)$ as considered in [41] (see also [2, Remark 5.3]).
@@
A strictly positive càdlàg process $D$ with $D_{0}\leq 1$ is a supermartingale deflator if $D(1+X)$ is a supermartingale, for every $X\in\mathcal{X}_{1}$. A strictly positive local martingale $D$ with $D_{0}=1$ is a local martingale deflator (LMD) if $DX^{(i,T)}$ is a local martingale, for every $(i,T)\in I_{0}\times\mathbb{R}_{+}$.
The next theorem is a version of the fundamental theorem of asset pricing based on the above notion of NUPBR. This result extends [18, Theorem A.1] in two directions: first, by considering an infinite time horizon; second, by proving the existence of a supermartingale deflator which is the reciprocal of a wealth process (numéraire portfolio, see [41]). Although the proof relies substantially on existing results such as [44, Theorem 1.7], we are not aware of a formulation that directly applies to our setting. We therefore provide a complete proof in Appendix A. We note that, in the present setup, the process $1+\widehat{X}$ appearing in Theorem 2.5 corresponds to the wealth process of the growth-optimal portfolio, a central object in the benchmark approach to finance (see, e.g., [51]).
@@
[…]
We point out that, at the present level of generality, NUPBR does not imply the existence of LMDs. Indeed, [18] gives an explicit example showing that in large financial markets market viability (in the sense of NUPBR, but also in the stronger sense of no asymptotic free lunch with vanishing risk) does only ensure the existence of supermartingale deflators, and not necessarily of LMDs. [^3] In the specific case of continuous-path processes NUPBR always implies the existence of an LMD, as shown in [45]. However, given a concrete model, it is usually difficult to characterize supermartingale deflators, since their definition is based on the set $\mathcal{X}_{1}$ and does not give information on the basic assets included in Definition 2.1. On the contrary, LMDs can be directly described in terms of the characteristics of the basic assets, as we are going to show in Section 3. Whenever it is nonempty, we denote by $\mathcal{D}$ the set of all LMDs.
@@ 3.1. Probabilistic setup
Let the filtered probability space $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ support a $d$-dimensional Brownian motion $W$ and an integer-valued random measure $\mu({\rm d}t,{\rm d}x)$ on $\mathbb{R}_{+}\times E$ with compensator $\nu({\rm d}t,{\rm d}x)=F_{t}({\rm d}x){\rm d}t$, where $E$ is a Polish space with its Borel sigma-field $\mathcal{B}(E)$ and $F_{t}({\rm d}x)$ is a kernel from $(\Omega\times\mathbb{R}_{+},\mathcal{P})$ into $(E,\mathcal{B}(E))$, with $\mathcal{P}$ denoting the predictable sigma-field on $\Omega\times\mathbb{R}_{+}$. The compensated random measure is denoted by $\tilde{\mu}({\rm d}t,{\rm d}x):=\mu({\rm d}t,{\rm d}x)-\nu({\rm d}t,{\rm d}x)$. We denote by $L^{2}_{\rm loc}(W)$ the set of all progressively measurable $\mathbb{R}^{d}$-valued processes $\theta=(\theta_{t})_{t\geq 0}$ such that $\int_{0}^{T}\|\theta_{t}\|^{2}{\rm d}t<+\infty$ a.s., for all $T>0$, and by $G_{\rm loc}(\mu)$ the set of all $\mathcal{P}\otimes\mathcal{B}(E)$-measurable functions $\varphi:\Omega\times\mathbb{R}_{+}\times E\to\mathbb{R}$ such that $\int_{0}^{T}\int_{E}((\varphi^{i}_{t}(x))^{2}\wedge|\varphi^{i}_{t}(x)|)F_{t}({\rm d}x){\rm d}t<+\infty$ a.s., for all $T>0$ (compare with [37, Theorem II.1.33-c]). We refer the reader to [37] for all unexplained notions of stochastic calculus.
@@
[…]
Assumption 3.1 implies that the integrals appearing in the forward rate equation (3.2) are well-defined for a.e. $T$. In addition, the integrability requirements appearing in parts (ii)-(iv) of Assumption 3.1 ensure the applicability of ordinary and stochastic Fubini theorems, in the versions of [59, Theorem 2.2] for the Brownian motion $W$ and [10, Proposition A.2] for the compensated random measure $\tilde{\mu}$. By [59, Remark 2.1], the mild measurability requirement in parts (ii)-(iii) holds if the processes $\alpha^{i}$ and $\beta^{i}$ are $\mathrm{Prog}\otimes\mathcal{B}(\mathbb{R}_{+})$-measurable, with $\mathrm{Prog}$ denoting the progressive sigma-field on $\Omega\times\mathbb{R}_{+}$.
@@
(2) Some recent works (see for instance [29, 30]) have considered HJM models that do not satisfy quasi-left-continuity (i.e., the set $\{(\omega,t)\in\Omega\times\mathbb{R}_{+}:\nu(\omega;\{t\}\times E)>0\}$ is not evanescent). Theorem 3.4 below can be generalized to this situation with an analogous proof. However, we do not pursue this generalization here since the SPDE analysis of Section 4 will necessitate quasi-left-continuity.
@@
$\leq\sum_{j=1}^{d}\int_{0}^{T}\left(\int_{0}^{u}(\beta^{i,j}(s,u))^{2}{\rm d}s\right)^{1/2}{\rm d}u<+\infty\qquad\text{a.s.},$
@@
$\leq T\int_{0}^{T}\int_{E}\int_{s}^{T}(\gamma^{i}(s,u,x))^{2}{\rm d}u\,\nu({\rm d}s,{\rm d}x)<+\infty\qquad\text{a.s.},$
thus ensuring that the stochastic integral $\int_{0}^{\cdot}\int_{E}\bar{\gamma}^{i}(s,x,T)\tilde{\mu}({\rm d}s,{\rm d}x)$ is well-defined as a local martingale, see [37, Theorem II.1.33]. The finiteness of the integral $\int_{0}^{\cdot}f^{i}(s,s){\rm d}s$ follows similarly as above under the validity of Assumption 3.1. An application of [37, Theorem II.8.10] to (3.3) gives the representation
@@
$\times\exp\left(\int_{0}^{t}f^{i}(s,s){\rm d}s-\int_{0}^{t}\bar{\alpha}^{i}(s,T){\rm d}s+\frac{1}{2}\int_{0}^{t}\|\bar{\beta}^{i}(s,T)\|^{2}{\rm d}s\right).$
The result of the lemma then follows by an application of Yor’s formula (see, e.g., [37, formula II.8.19]), making use of (3.4) and recalling that $S^{i}=S^{i}_{0}\mathcal{E}(Z^{i})$, where $Z^{i}$ is given by (3.1). ∎